# A chaotic lattice field theory in two dimensions

### v2

@@ Abstract
A temporally chaotic system is exponentially unstable with time: double the time, and exponentially more orbits are required to cover its strange attractor to the same accuracy. For a system of large spatial extent, the complexity of the spatial shapes also needs to be taken into account; double the spatial extent, and exponentially as many distinct spatial patterns might be required to describe the repertoire of spatial shapes to the same accuracy.  [1, 2, 3] Based on the insight that temporal and spatial instabilities can be treated on equal footing, the ‘chronotopic’ theory of 1990’s offers an elegant description of such spatiotemporal chaos, with predictions of the theory encapsulated in an ‘entropy density’ function. [4, 5, 6, 7] The first decade of 2000’s saw rapid progress in description of transitional turbulence in fluids, in terms of ‘exact coherent structures’, numerically exact spatiotemporally periodic unstable solutions of Navier-Stokes equations. [8, 9, 10, 11, 12, 13, 14] This approach to turbulence aims to predict measurable observables from the defining (Navier-Stokes) equations, without any statistical assumptions. But how are these solutions to be pieced together? [15] As a single such solution requires highly unstable time integration over $10^{4}-10^{6}$ discretization ODEs, estimates of ‘chronotopic’ theory observables based on large spacetime volume limits of numerical simulations are out of question. In this paper we develop a new spatiotemporal theory of turbulence, which assigns a new kind of a global weight, an ‘orbit Jacobian’, to each exact coherent structure, with a new enumeration of global solutions as sums over all Bravais lattices. Our deterministic field theory replaces numerical extrapolations of finite spacetime simulations of the 1990’s chronotopy by our main result, the exact spatiotemporal deterministic zeta function, with computable cycle expansion truncations errors, decreasing exponentially with the spatiotemporal volumes of periodic states included in its evaluation.
@@
Our goal here is to make this ‘spatiotemporal chaos’ tangible and precise (see Sec. X.1), in a series of papers that introduce its theory and its implementations. The companion paper I [16] focuses on the $1d$ chaotic lattice field theory, and a novel treatment of time-reversal invariance. In this paper, paper II, we develop the theory of 2dd spatiotemporal chaotic systems; and in the companion paper III [17] we apply the theory to several nonlinear field theories. As our intended audience spans many disjoint specialties, from fluid dynamics to quantum field theory, the exposition entails much pedagogical detail, so let us start by stating succinctly what the key novelty of our theory is.
@@
There are two ways of studying stabilities of translationally-invariant systems, illustrated by Fig. 8 (b) and (c):
@@
Much work follows, [20] with some details elaborated in Sec. VIII.
(ii) A crystallographer, or field theorist starts with an infinite lattice, or continuum spacetime. The approach –as we show here in Sec. IX.2– yields weights that are multiplicative for repeats of spatiotemporally periodic solutions,
@@
(quotation marks, as the precise statement is in terms of stability exponents rather than determinants). This fact simplifies everything, and yields the main result of this paper, the spatiotemporal zeta function (Sec. X.4), expressed as a product of Euler functions $\phi(t_{p})$, one for each prime spatiotemporal solution of weight $t_{p}$,
@@
In our formulation the ‘chaos theory’ is a Euclidean deterministic field theory. But, as developed here, this theory looks nothing like the textbook expositions [26, 27, 28, 29, 30, 31, 20] of temporally chaotic, few degrees-of-freedom dynamical systems. There one is given an initial state, which then evolves in time, much like in mechanics, where given an initial phase-space point, the integration of Hamilton’s equations traces out a phase-space trajectory. For a reader versed in fluid dynamics or atomic physics the most disconcerting aspect of the field-theoretic perspective is that time is just one of the coordinates over which a field configuration is defined: each field-theoretic solution is a static solution over the infinite spacetime. There is no ‘evolution in time’, no stable / unstable manifolds, no ‘hyperbolicity’, no ‘mixing’.
Furthermore –just as the discretization of time by Poincaré sections aids analysis of temporal chaos– we find it convenient to discretize both time and space. Spatiotemporally steady turbulent flows offer one physical motivation for considering such models: a rough approximation to such flows is obtained discretizing them into spatiotemporal cells, with each cell turbulent, and cells coupled to their neighbors. Lattices also arise naturally in many-body problems, such as many-body quantum chaos models studied in Refs. 32, 33, 34, 35, see Sec. III.2.
Outline. We start our formulation of chaotic field theory (Sec. I) by defining the field theory partition sums in terms of spatiotemporally periodic states (Sec. I.1). A deterministic field theory has support on the set of all solutions of systems’ defining equations. Its building blocks are periodic states, spatiotemporally periodic solutions of system’s defining equations (Sec. II).
In Sec. III we introduce the field theories studied here. In particular, the simplest of chaotic field theories, the spatiotemporal cat [36, 37] of Sec. III.1, a discretization of the compact boson Klein-Gordon equation,
@@
In Sec. III.2 we motivate of choice of nonlinear field theories studied here by the 1955 Fermi-Pasta-Ulam-Tsingou [43, 44] model which illustrates how many-body ‘chaos theory’ morphs into a Euclidian strong-coupling, anti-integrable deterministic field theory.
@@
Crucial to ‘chaos’ is the notion of stability: in Sec. IV we describe spatiotemporal stability of above field theories’ periodic states in terms of their orbit Jacobian operators. Periodic orbit theory for a time-evolving dynamical system on a one-dimensional temporal lattice is organized by grouping orbits of the same time period together. [18, 19, 45, 20, 16] For systems characterized by several translational symmetries, one has to take care of multiple periodicities; in the language of crystallography, organize the periodic orbit sums by corresponding Bravais lattices, or, in the language of field theory, by the ‘sum over geometries’. In Sec. V we enumerate and construct spacetime geometries, or $d=2$ Bravais lattices ${L}\!\times\!{T}]_{S}$}, of increasing spacetime periodicities. The classification of periodic states proceeds in two steps. On the coordinate level, periodicity is imposed by the hierarchy of Bravais lattices of increasing periodicities. On the field-configuration level, the key to the spatiotemporal periodic orbit theory is the enumeration and determination of prime orbits, the basic building blocks of periodic orbit theory (Sec. VI).
@@
The central idea of spatiotemporal theory developed here, global orbit stability, has its origins in the 1886 work of Hill [46] and Poincaré. [47] The likelihood of each solution is given by the orbit Jacobian, the determinant of its spatiotemporal orbit Jacobian matrix. Compared to the temporal-evolution chaos theory, the orbit Jacobian (more precisely, the stability exponent) is the central innovation of our field-theoretic formulation of chaotic field theory, so we return to it throughout the paper. The calculations are carried out on the reciprocal lattice (Sec. VII). We discuss primitive cell computations in Sec. VIII, as prequel to introducing the stability exponent of a periodic state over spatiotemporally infinite Bravais lattice, computed on the reciprocal lattice (Sec. IX). For spatiotemporal cat we evaluate and cross-check orbit Jacobians by two methods, either on the reciprocal lattice (Sec. VIII.2), or by the ‘fundamental fact’ evaluation (Appendix B.3).
@@
Having enumerated all Bravais lattices (Sec. V), determined periodic states over each (Sec. VI.1), computed the weight of each periodic state (Sec. IX), we can now write down the deterministic field theory partition sum as a sum over all spatiotemporal solutions of the theory (Sec. X). In Sec. X.3 we reexpress the partition sum in terms of prime orbits, and in Sec. X.4 we construct the spatiotemporal zeta function. What makes these resummations possible is the multiplicative property of orbit Jacobians announced in Eq. (2), provided by their evaluation over the spatiotemporally infinite Bravais lattice (Sec. IX), the key property that is violated by finite-dimensional matrix approximations that are the basis of the traditional Gutzwiller-Ruelle temporal periodic orbits theory (Sec. VIII). In Sec. X.6 explain how one computes expectation values of observables in deterministic chaotic field theories.
@@
Finally, we know that time-evolution cycle-expansions’ convergence is accelerated by shadowing of long orbits by shorter periodic orbits. [49] In Sec. XI we check numerically that spatiotemporal cat periodic states that share finite spatiotemporal mosaics shadow each other to exponential precision. We presume (but do not show) that this shadowing property ensures that the predictions of the theory are dominated by the shortest period prime orbits.
This completes our generalization [37, 16, 50, 17] of the temporal-evolution deterministic chaos theory [20, 45] to spatiotemporal chaos / turbulence, and recasts both in the formalism of conventional solid state physics, field theory, and statistical mechanics. Our results are summarized and open problems discussed in Sec. XII. Our calculations are reported in Appendices. Icon [Uncaptioned image] on the margin links a block of text to a supplementary online video. For additional material, online talks and related papers, see ChaosBook.org/overheads/spatiotemporal.
I Lattice field theory
@@
spanned by a set of orthogonal vectors $\mathbf{e}_{j}$, with lattice spacing $a_{j}=|\mathbf{e}_{j}|=\Delta x_{j}$ along the direction of vector $\mathbf{e}_{j}$. We shall use lattice units, almost always setting $a_{j}=1$ (for another, modular function parametrization choice, see Eq. (75)). A field $\phi(x)$ over $d$ continuous coordinates $x_{j}$ is represented by a discrete array of field values over lattice sites
@@
While we refer here to such discretization as a lattice field theory, the lattice might arise naturally from a many-body setting with the nearest neighbors interactions, such as many-body quantum chaos models studied in Refs. 32, 33, 34, 35, with a multiplet of fields at every site. [36]
I.1 Periodic field configurations
@@
where the $d\!\times\!d]$ basis matrix $\mathbb{A}=\left[\mathbf{a}_{1},\mathbf{a}_{2},\cdots,\mathbf{a}_{d}\right]$ formed from primitive integer lattice vectors $\{\mathbf{a}_{j}\}$ defines a $d$-dimensional primitive cell[53, 54] If the lattice spacing Eq. (6) is set to 1, the lattice volume, or the volume of primitive cell
@@
tile periodically the doubly-infinite field configuration Eq. (7).
We focus on the one-dimensional case for now, postpone discussion of higher-dimensional Bravais lattices to Sec. V.
I.2 Orbits
@@
While each field configuration $\mathsf{r}^{j}\mathsf{\Phi}$ might be a distinct point in the primitive cell’s state space Eq. (12), they are equivalent, in the sense that they all are the same set of lattice site fields $\{{\phi}_{z}\}$, up to a cyclic relabelling of lattice sites.
@@
each a set of equivalent field configurations, labelled perhaps by $\mathsf{\Phi}_{p}$, one of the configurations in the set. Each orbit is a fixed point of $T$, as for any translation $\mathsf{r}^{j}\mathcal{M}_{p}=\mathcal{M}_{p}\,.$ The number of distinct field configurations in the orbit is $n_{p}$, the period of the orbit. (For orbits over two-dimensional lattices, see Sec. VI).
I.3 Prime orbits over one-dimensional primitive cells
@@
The distinction arises in enumeration of field configurations over a primitive cell. The primitive cell state spaces Eq. (12) are here 6-, 3-dimensional, respectively. Both orbits depend on the same 3 distinct lattice site field values $\phi_{z}$. On the primitive cell $2\mathbb{A}$, however, the 6 lattice sites field configuration $\mathsf{\Phi}_{2\mathbb{A}}$ is not prime, it is a repeat of the field configuration $\mathsf{\Phi}_{\mathbb{A}}$. We elaborate this distinction in Sec. IV.2.
This is how ‘prime periodic orbits’ and their repeats work for the one-dimensional, temporal lattice. For a two-dimensional square lattice the notion of ‘prime’ is a bit trickier, so we postpone its discussion to Sec. VI.

### v3

@@ Abstract
A temporally chaotic system is exponentially unstable with time: double the time, and exponentially more orbits are required to cover its strange attractor to the same accuracy. For a system of large spatial extent, the complexity of the spatial shapes also needs to be taken into account; double the spatial extent, and exponentially as many distinct spatial patterns might be required to describe the repertoire of spatial shapes to the same accuracy.  [1, 2, 3] Based on the insight that temporal and spatial instabilities can be treated on equal footing, the ‘chronotopic’ theory of 1990s offers an elegant description of such spatiotemporal chaos, with predictions of the theory encapsulated in an ‘entropy density’ function. [4, 5, 6, 7] The first decade of 2000s saw rapid progress in description of transitional turbulence in fluids, in terms of ‘exact coherent structures’, numerically exact spatiotemporally periodic unstable solutions of Navier-Stokes equations. [8, 9, 10, 11, 12, 13, 14] This approach to turbulence aims to predict measurable observables from the defining (Navier-Stokes) equations, without any statistical assumptions. But how are these solutions to be pieced together? [15] As a single such solution requires highly unstable time integration over $10^{4}-10^{6}$ discretization ODEs, estimates of ‘chronotopic’ theory observables based on large spacetime volume limits of numerical simulations are out of question. In this paper we develop a new spatiotemporal theory of turbulence, which assigns a new kind of a global weight, an ‘orbit Jacobian’, to each exact coherent structure, with a new enumeration of global solutions as sums over all Bravais lattices. Our deterministic field theory replaces numerical extrapolations of finite spacetime simulations of the 1990s chronotopy by our main result, the exact spatiotemporal deterministic zeta function, with computable cycle expansion truncations errors, decreasing exponentially with the spatiotemporal volumes of periodic states included in its evaluation.
I Introduction
@@
Our goal here is to make this ‘spatiotemporal chaos’ tangible and precise (see Sec. XI.1), in a series of papers that introduce its theory and its implementations. The companion paper I [16] focuses on the $1d$ chaotic lattice field theory, and a novel treatment of time-reversal invariance. In this paper, paper II, we develop the theory of 2dd spatiotemporal chaotic systems; and in the companion paper III [17] we apply the theory to several nonlinear field theories. As our intended audience spans many disjoint specialties, from fluid dynamics to quantum field theory, the exposition entails much pedagogical detail, so let us start by stating succinctly what the key novelty of our theory is.
$(a)$    Refer to caption
$(b)$    Refer to caption
$(c)$    Refer to caption
Figure 1: A one-dimensional temporal lattice period-5 periodic state $\mathsf{\Phi}_{c}=[{\phi}_{0}\,{\phi}_{1}\,{\phi}_{2}\,{\phi}_{3}\,{\phi}_{4}]$ illustrated by (a) five repeats of the primitive cell periodic state. (b) An internal perturbation $h_{z}$ has the same periodicity as the periodic state. Its spectrum, evaluated in Sec. IX, is discrete. (c) A transverse perturbation $h_{z}$ is an arbitrary, aperiodic function over the infinite lattice. Its spectrum, evaluated by the Floquet-Bloch theorem in Sec. X, is a continuous function of wave number $k$. Horizontal: lattice sites labeled by $z\in\mathbb{Z}$. Vertical: (a) value of field $\phi_{z}$, (b-c) perturbation $h_{z}$, plotted as a bar centered at lattice site $z$. Values of the field and perturbation are shown in blue within the primitive cell, and in orange outside the primitive cell.
@@
There are two ways of studying stabilities of translationally invariant systems, illustrated by Fig. 1 (b) and (c):
@@
Much work follows, [20] with some details elaborated in Sec. IX.
(ii) A crystallographer, or field theorist starts with an infinite lattice, or continuum spacetime. The approach –as we show here in Sec. X.2– yields weights that are multiplicative for repeats of spatiotemporally periodic solutions,
@@
(quotation marks, as the precise statement is in terms of stability exponents rather than determinants). This fact simplifies everything, and yields the main result of this paper, the spatiotemporal zeta function (Sec. XI.4), expressed as a product of Euler functions $\phi(t_{p})$, one for each prime spatiotemporal solution of weight $t_{p}$,
@@
In our formulation the ‘chaos theory’ is a Euclidean deterministic field theory. However, as developed here, this theory looks nothing like the textbook expositions [26, 27, 28, 29, 30, 31, 20] of temporally chaotic, few degrees-of-freedom dynamical systems. There one is given an initial state, which then evolves in time, much like in mechanics, where given an initial phase-space point, the integration of Hamilton’s equations traces out a phase-space trajectory. For a reader versed in fluid dynamics or atomic physics the most disconcerting aspect of the field-theoretic perspective is that time is just one of the coordinates over which a field configuration is defined: each field-theoretic solution is a static solution over the infinite spacetime. There is no ‘evolution in time’, no stable / unstable manifolds, no ‘hyperbolicity’, no ‘mixing’.
Furthermore –just as the discretization of time by Poincaré sections aids analysis of temporal chaos– we find it convenient to discretize both time and space. Spatiotemporally steady turbulent flows offer one physical motivation for considering such models: a rough approximation to such flows is obtained discretizing them into spatiotemporal cells, with each cell turbulent, and cells coupled to their neighbors. Lattices also arise naturally in many-body problems, such as many-body quantum chaos models studied in Refs. 32, 33, 34, 35, see Sec. IV.2.
Outline. We start our formulation of chaotic field theory (Sec. II) by defining the field theory partition sums in terms of spatiotemporally periodic states (Sec. II.1). A deterministic field theory has support on the set of all solutions of systems’ defining equations. Its building blocks are periodic states, spatiotemporally periodic solutions of system’s defining equations (Sec. III).
In Sec. IV we introduce the field theories studied here. In particular, the simplest of chaotic field theories, the spatiotemporal cat [36, 37] of Sec. IV.1, a discretization of the compact boson Klein-Gordon equation,
@@
In Sec. IV.2 we motivate of choice of nonlinear field theories studied here by the 1955 Fermi-Pasta-Ulam-Tsingou [43, 44] model which illustrates how many-body ‘chaos theory’ morphs into a Euclidian strong-coupling, anti-integrable deterministic field theory.
@@
Crucial to ‘chaos’ is the notion of stability: in Sec. V we describe spatiotemporal stability of above field theories’ periodic states in terms of their orbit Jacobian operators. Periodic orbit theory for a time-evolving dynamical system on a one-dimensional temporal lattice is organized by grouping orbits of the same time period together. [18, 19, 45, 20, 16] For systems characterized by several translational symmetries, one has to take care of multiple periodicities; in the language of crystallography, organize the periodic orbit sums by corresponding Bravais lattices, or, in the language of field theory, by the ‘sum over geometries’. In Sec. VI we enumerate and construct spacetime geometries, or $d=2$ Bravais lattices ${L}\!\times\!{T}]_{S}$}, of increasing spacetime periodicities. The classification of periodic states proceeds in two steps. On the coordinate level, periodicity is imposed by the hierarchy of Bravais lattices of increasing periodicities. On the field-configuration level, the key to the spatiotemporal periodic orbit theory is the enumeration and determination of prime orbits, the basic building blocks of periodic orbit theory (Sec. VII).
@@
The central idea of spatiotemporal theory developed here, global orbit stability, has its origins in the 1886 work of Hill [46] and Poincaré. [47] The likelihood of each solution is given by the orbit Jacobian, the determinant of its spatiotemporal orbit Jacobian matrix. Compared to the temporal-evolution chaos theory, the orbit Jacobian (more precisely, the stability exponent) is the central innovation of our field-theoretic formulation of chaotic field theory, so we return to it throughout the paper. The calculations are carried out on the reciprocal lattice (Sec. VIII). We discuss primitive cell computations in Sec. IX, as prequel to introducing the stability exponent of a periodic state over spatiotemporally infinite Bravais lattice, computed on the reciprocal lattice (Sec. X). For spatiotemporal cat we evaluate and cross-check orbit Jacobians by two methods, either on the reciprocal lattice (Sec. IX.2), or by the ‘fundamental fact’ evaluation (Appendix B.3).
@@
Having enumerated all Bravais lattices (Sec. VI), determined periodic states over each (Sec. VII.1), computed the weight of each periodic state (Sec. X), we can now write down the deterministic field theory partition sum as a sum over all spatiotemporal solutions of the theory (Sec. XI). In Sec. XI.3 we reexpress the partition sum in terms of prime orbits, and in Sec. XI.4 we construct the spatiotemporal zeta function. What makes these resummations possible is the multiplicative property of orbit Jacobians announced in Eq. (2), provided by their evaluation over the spatiotemporally infinite Bravais lattice (Sec. X), the key property that is violated by finite-dimensional matrix approximations that are the basis of the traditional Gutzwiller-Ruelle temporal periodic orbits theory (Sec. IX). In Sec. XI.6, we explain how one computes expectation values of observables in deterministic chaotic field theories.
@@
Finally, we know that time-evolution cycle-expansions’ convergence is accelerated by shadowing of long orbits by shorter periodic orbits. [49] In Sec. XII we check numerically that spatiotemporal cat periodic states that share finite spatiotemporal mosaics shadow each other to exponential precision. We presume (but do not show) that this shadowing property ensures that the predictions of the theory are dominated by the shortest period prime orbits.
This completes our generalization [37, 16, 50, 17] of the temporal-evolution deterministic chaos theory [20, 45] to spatiotemporal chaos / turbulence, and recasts both in the formalism of conventional solid state physics, field theory, and statistical mechanics. Our results are summarized and open problems discussed in Sec. XIII. Our calculations are reported in Appendices. Icon [Uncaptioned image] on the margin links a block of text to a supplementary online video. For additional material, online talks and related papers, see ChaosBook.org/overheads/spatiotemporal[51]
II Lattice field theory
@@
spanned by a set of orthogonal vectors $\mathbf{e}_{j}$, with lattice spacing $a_{j}=|\mathbf{e}_{j}|=\Delta x_{j}$ along the direction of vector $\mathbf{e}_{j}$. We shall use lattice units, almost always setting $a_{j}=1$ (for another, modular function parameterization choice, see Eq. (75)). A field $\phi(x)$ over $d$ continuous coordinates $x_{j}$ is represented by a discrete array of field values over lattice sites
@@
While we refer here to such discretization as a lattice field theory, the lattice might arise naturally from a many-body setting with the nearest neighbor interactions, such as many-body quantum chaos models studied in Refs. 32, 33, 34, 35, with a multiplet of fields at every site. [36]
II.1 Periodic field configurations
@@
where the $d\!\times\!d]$ basis matrix $\mathbb{A}=\left[\mathbf{a}_{1},\mathbf{a}_{2},\cdots,\mathbf{a}_{d}\right]$ formed from primitive integer lattice vectors $\{\mathbf{a}_{j}\}$ defines a $d$-dimensional primitive cell[54, 55] If the lattice spacing Eq. (6) is set to 1, the lattice volume, or the volume of the primitive cell
@@
Figure 2: The intersections of the light gray lines -lattice sites $z\in\mathbb{Z}^{2}$- form the integer square lattice Eq. (6). (a) Translations of the primitive cell $\mathbb{A}={[3\!\times\!2]_{1}}$ spanned by primitive vectors $\mathbf{a}_{1}=(3,0)$ and $\mathbf{a}_{2}=(1,2)$ define the Bravais lattice $\mathcal{L}_{\mathbb{A}}$. (b) The primitive vectors $\mathbf{a}_{1}=(2,-2)$ and $\mathbf{a}_{2}=(-1,4)$ form a primitive cell $\mathbb{A}^{\prime}$} equivalent to (a) by a unimodular transformation. The intersections (red points) of either set of dashed lines form the same Bravais lattice $\mathcal{L}_{\mathbb{A}}=\mathcal{L}_{\mathbb{A}^{\prime}}$. The volume Eq. (11) of either primitive cell is ${N}_{\mathcal{L}}=6$, the number of integer lattice sites within the cell, with the tips of primitive vectors and tiles’ outer boundaries belonging to the neighboring tiles. This will be discussed further in Fig. 6.
@@
tile periodically the doubly infinite field configuration Eq. (7).
We focus on the one-dimensional case for now, postpone discussion of higher-dimensional Bravais lattices to Sec. VI.
II.2 Orbits
@@
While each field configuration $\mathsf{r}^{j}\mathsf{\Phi}$ might be a distinct point in the primitive cell’s state space Eq. (12), they are equivalent, in the sense that they all are the same set of lattice site fields $\{{\phi}_{z}\}$, up to a cyclic relabeling of lattice sites.
@@
each a set of equivalent field configurations, labeled perhaps by $\mathsf{\Phi}_{p}$, one of the configurations in the set. Each orbit is a fixed point of $T$, as for any translation $\mathsf{r}^{j}\mathcal{M}_{p}=\mathcal{M}_{p}\,.$ The number of distinct field configurations in the orbit is $n_{p}$, the period of the orbit. (For orbits over two-dimensional lattices, see Sec. VII).
II.3 Prime orbits over one-dimensional primitive cells
@@
The distinction arises in enumeration of field configurations over a primitive cell. The primitive cell state spaces Eq. (12) are here 6-, 3-dimensional, respectively. Both orbits depend on the same three distinct lattice site field values $\phi_{z}$. On the primitive cell $2\mathbb{A}$, however, the six lattice sites field configuration $\mathsf{\Phi}_{2\mathbb{A}}$ is not prime, it is a repeat of the field configuration $\mathsf{\Phi}_{\mathbb{A}}$. We elaborate this distinction in Sec. V.2.
This is how ‘prime periodic orbits’ and their repeats work for the one-dimensional, temporal lattice. For a two-dimensional square lattice the notion of ‘prime’ is a bit trickier, so we postpone its discussion to Sec. VII.